Two alkenes, A and B, each have the formula C5H10. Both react with H2/Pt and with

HBr to give identical products. What are the structures of A and B?


1
Expert's answer
2021-02-26T07:43:23-0500

The two alkenes are pentene isomers with structures:

1. H3C-CH2-CH=CH-CH3

2. H3C-CH=CH-CH2-CH3